# Refusal Tokens: A Simple Way to Calibrate Refusals in Large Language Models

### v1

@@ 5 Test-Time Control Using [Refuse] and [Respond] Tokens
The refusal token introduces test-time capabilities. By training with the refusal token, the refusal rate can be altered at test-time. This ability cannot occur when training without the token. The model predicts this token, providing a softmax probability associated with the refusal token. This token probability can be interpreted as the confidence with which the model “thinks” the question should respond with a refusal message. Conversely, the response token is interpreted as the probability that the model should respond. As this probability may not be perfectly calibrated, we sweep different thresholds to find different refusal rates. We generate the token if $p(\text{{[refuse]}}|C(x))>T$, where $T$ is a threshold set by the user. By adjusting the threshold, $T$, we demonstrate that the refusal rates can be effectively controlled.
Refusal tokens provides control of the refusal rate. We sweep the thresholds of the refusal token across the two settings–training with and without contrast training examples–to observe the change in the true positive and false positive rates. In Figure 2, the threshold provides control over the true positive and false positive rates. Figures 2(a) and 2(b) show that adding contrast data (SFT data that lies close the boundary between the two classes but are non-refusal) results in a better Pareto frontier than training without the token.
(a) CoCoNot with no contrast in training data
@@
Figure 2: Manipulating the refusal token provides different refusal rates at test-time without retraining. The left and right figures show that both true positive and false positive rates on CoCoNot eval change as we vary the threshold of the refusal token. The models are trained with ultrachat and refusal messages from the CoCoNot training data. Left is trained without any contrast data, and the right is trained with contrast data, which is one-tenth of the refusal data.
@@ 5.1 Controlling Individual Types of Instructions with Category Refusal Tokens
We now experiment with having five distinct refusal tokens that differentiate between refusal types for CoCoNot. Additionally, we consider the temporal setting with one temporal refusal token. For all experiments in this section, we add refusals and/or contrast data to UltraChat.
@@
Thresholding schemes and logit bias. We explore two types of thresholding strategies: (1) category thresholding, refusing with that category token if a token from selected category tokens is the highest probability among the refusal tokens and rises above a threshold, and (2) sum thresholding, refusing only if the sum of all category token scores exceeds a threshold. For category thresholding, we emit the refusal token that is the highest probability among the refusal tokens and is in the selected category tokens; otherwise, we emit the token with the highest probability. For sum thresholding, we emit the category refusal token highest probability when the condition described earlier is met; otherwise, we emit a response token. Algorithmic versions of these schemes can be found in Appendix A.6. For logit bias, we manipulate the sensitivity of different refusal types by adding a constant bias to the (unnormalized) logits for the refusal tokens.
@@
Figure 3: Individual category refusal tokens enable precise control over query types. Refusal rates for different categories on CoCoNot when category-specific tokens are suppressed or not generated by the model. The blue dashed bars suppress the refusal token in a model trained exclusively with response and refusal tokens, no category tokens. By suppressing tokens from specific categories during inference, we demonstrate control over the types of refusals. The two dashed bars per group reflect the effect of suppressing a category’s token, either through category-specific suppression or a single refusal token. We also observe category overlap with both these experiments and a manual inspection; for instance, Humanizing Requests may fall into multiple categories.
@@
Independent control of sensitivity for different refusal types. To test whether categories can be independently controlled, we completely suppress each token one-at-a-time, and observe the impact of this suppression on other (non-suppressed) refusal types. In Figure 3, we observe that the sensitivity of each refusal category can be adjusted with little impact on other categories of refusals. There is an exception though: Humanizing Requests proved particularly difficult to suppress and did not respond to their token as other categories did. After inspecting the questions and responses of the Humanizing Requests category, we found that many of the questions contained questions or instructions similar to other categories.
@@
Thus, many of the Humanizing questions or instructions are classified as one of the other refusal categories, i.e. the model emitted the incorrect refusal token. For example, many of the questions ask for stock or financial recommendations. These types of requests could easily be refused due to temporal issues (no access to real-time information), input modality issues (needing access to current portfolios), or safety (not wanting to provide financial information). Nevertheless, Figure 3 highlights that one can use individual category tokens to control individual distributions.
@@
Figure 4: Thresholding the refusal tokens increase F1 scores and controls the true positive and false positive rates for a single instruction type (temporal setting). For our temporal experimental setting, we train UltraChat with $2$k refusals and $2$k contrast examples. The left shows thresholding achieves a better F1 Score, and the right shows thresholding controls the true positive false positive rates.
@@
Increasing F1 scores via category refusal tokens. We first consider our temporal setting. Particularly, we sweep the thresholds of a model trained with UltraChat, $\sim 2$k temporal refusal messages, and $\sim 2$k temporal contrast training examples. We experiment with values of $T$ from 00 to $1$ in increments of $0.1$, where we only sweep one token. In Figure 4, we observe that F1 scores improve when properly calibrating the thresholds, finding that $T=0.1$ performs the best. It is worth noting, through experiments, we found training solely with the underlying SFT dataset, training without additional refusal or contrast data, leaves the model with an inherent refusal rate. In Figure 4, the false positive rate does not drop below approximately $0.35$.
@@
Table 2: Using category-wise thresholding and logit bias to increase the refusal rates of particular categories, a case study. We apply the category-wise threshold at $T=0.1$ or a logit bias of $B=4$ to two queries types simultaneously: Humanizing and Indeterminate. This experiment shows that manipulating a subset of categories increases overall F1 performance without retraining the model. In contrast, thresholding a refusal single token yields higher refusal rates across all categories, notably, doubling the contrast refusal rate. The numbers on the left side of the vertical line are the rates that we expect to change by thresholding or logit bias.
@@
Moreover, we confirm the chosen thresholds are ideal. We perform a “cheap sweep” where we utilize only the tokens for the labels instead of the llm-as-a-judge model. More concretely, for each query we apply a single forward pass to find which token is emitted and use this token as the label. This allows us to find thresholds without generating full responses to the queries. We find the best refusal thresholds by independently finding the threshold that maximizes the F1 Score for each category token from $0.1$ to $0.9$ in increments of $0.1$. From this, we found that both Humazing and Interdetemine categories had the best threshold at $0.1$. Thus, confirming the thresholds chosen in our case study is ideal. This “cheap sweep” can be utilized to further improve F1 scores across all categories.
@@
In Table 2, using category-wise thresholding and logit bias, the refusal rates increased for Humanizing by $\sim 5\%$ for both thresholding and logit bias and Interdetermined by $8.0\%$ for thresholding and $5.2\%$ for logit bias. These test-time approaches improved the F1 score. Conversely, when setting the single token to a threshold of $T=0.1$, the contrast refusal rate (Type II error) doubles, increasing refusal rates in all categories. Thus, individually controlling the different category-wise refusal tokens at test-time leads to more control on category refusal rates, whether utilizing either category-wise thresholding or logit bias.
Improving F1 scores with sum thresholding. The sum thresholding scheme is considered where controlling individual categories is not of interest. Particularly, we sweep the thresholds of a model trained with UltraChat, CoCoNot refusal messages, and CoCoNot contrast training examples. In Figure 5, by sweeping the thresholds between 00 and $1$ in increments of $0.1$, a threshold of $0.6$ yields the best F1 score over sampling. This experiment further shows that category tokens can be altered in different ways at test-time for better F1 performance or different needs. Overall, using multiple tokens provides greater flexibility and steerability for the user than a single refusal token. However, if a user does not require this level of flexibility or prefers not to add many new tokens to the vocabulary, a single token remains an excellent solution for controlling the model’s refusal rate, as shown in Figure 2. Ultimately, the choice depends on the user’s specific preferences and requirements.
Figure 5: Sum thresholding is another way to effectively utilize the category tokens at test-time. (Left) F1 scores on CoCoNot evaluation, (center) average of the refusal rates for refusal categories in the CoCoNot evaluation, and (right) the refusal rate the contrast category in the CoCoNot evaluation as the threshold is swept. The refusal token is emitted if the sum of the scores for all category tokens exceeds the threshold. At a threshold of $T=0.6$, the F1 Score is highest at $0.946$ up from $0.938$, cutting the error rate by $\sim 12\%$.
6 Out-of-the-Box Benefits
A major advantage of incorporating refusal tokens lies in their ability to influence model behavior at test-time. Notably, even without using the refusal tokens to control a model at test-time, the mere inclusion of refusal tokens during training enhances the model’s refusal behaviors (measured by F1 scores). In our primary experimental setup, we focus on training with temporal refusals and/or temporal contrast data, as outlined in Section 4. These experiments examine how fine-tuning a model on refusal data from one type of query affects the refusal rates for other types of questions. Additionally, we assess how introducing the refusal token influences the refusal behavior, without applying test-time interventions.
We begin by evaluating a model trained with the Alpaca dataset, including only temporal refusal data (i.e., excluding contrast training data), to observe its impact on Type I and Type II errors. Moreover, we explore how the refusal token itself shapes refusal behavior, particularly concerning these errors. To better understand the relationship between the quantity of refusal data and the model’s refusal rates, we experiment with varying proportions of $2$k refusal examples–$1\%,5\%,10\%,50\%,100\%$–integrated into the Alpaca dataset. This range allows us to analyze how different amounts of refusal data influence the model’s refusal performance across question types, beyond what is explicitly represented in the training set.
From Figure 6 (left), very few refusal messages in the training data are required for other types of refusal questions to be affected. Particularly, with only $200$ refusal messages, CoCoNot queries and TriviaQA questions refusal rate increase. Thus, this highlights a model trained to refuse specific instruction types will refuse other instruction types without explicitly training to refuse those queries. Furthermore, from Figure 6 (left), the refusal token can limit this Type II error, but as you scale the number of examples, this benefit is limited.
Figure 6: The token limits Type II error in an out-of-the-box setting but is not sufficient as the refusal rate increases across the board. Left are refusal rates on the three subsets of the evaluation: temporal questions, CoCoNot questions, and TriviaQA questions, where one model is trained with the refusal token and one without the token. Right are refusal rates on the three subsets of the evaluation: temporal questions, CoCoNot questions, and TriviaQA questions where one model is trained with contrast data and one without contrast data, where both models are trained with the refusal token. The x-axis is how many instructions the model was trained with. The gray line represents the rates with no refusal messages in the instruction data.
As data is the key to LLM training, we add contrast data to understand how adding borderline examples affects the refusal rates. In our experiments, we add one contrast instruction with one refusal instruction in SFT training data, adding the refusal token to all experiments. From Figure 6 (right), adding the contrast data to the training dataset limits the refusal rates on other instruction types as the number refusals scales. Thus, in situations where you only want to refuse a particular instruction type, i.e. limit Type II error, including contrast data in the training data is crutical.
Furthermore, we explore the case where the balance of contrast to refusal messages is one to ten, which is the case for CoCoNot training dataset. In Table 3, even when training with this imbalance the contrast training data limits the amount the refusal rate on innocuous questions, albeit not to the same refusal rates as not training with refusals. Additionally, from the table, adding both a single refusal token and category tokens improves F1 scores under default sampling methods. However, we suspect the exact benefits might be model and hyperparameter dependent. Nevertheless, we see benefits in all models that we explored (llama-3.1 and Mistral (Jiang et al., 2023)) in Table 7 in the Appendix.
@@ 6 Out-of-the-Box Benefits
Table 3: Refusal tokens and contrast data improve F1 performance on CoCoNot without thresholding at test-time. Ablation studies on training with CoCoNot refusal messages, refusal tokens, and contrast data. We evaluate llama-3 $8$B performance across different tasks including MMLU (Hendrycks et al., 2020), ARC tasks (Clark et al., 2018), HellaSwag (Zellers et al., 2019), and TruthfulQA MC2 (Lin et al., 2022), following hyperparameters from Tunstall et al. (2023).
@@
7 Discussion

### v2

@@ 5 Test-Time Control Using [Refuse] and [Respond] Tokens
(a) CoCoNot w/o contrast in training data
Figure 2: Manipulating the refusal token provides different refusal rates at test time without retraining. The left and right figures show that both true positive and false positive rates on CoCoNot evaluation change as we vary the threshold of the refusal token. The models are trained with UltraChat and refusal messages from the CoCoNot training data. The left model is trained with contrast data, which constitutes one-tenth of the refusal data size, and the right is trained without any contrast data.
The refusal token introduces test-time capabilities. By training with the refusal token, the refusal rate can be altered at test-time. The model predicts this token, providing a softmax probability associated with the refusal token. This token probability can be interpreted as the confidence with which the model “thinks” the question should respond with a refusal message. Conversely, the response token is interpreted as the probability that the model should respond. As this probability may not be perfectly calibrated, we sweep different thresholds to find different refusal rates. We generate the token if $p(\text{{[refuse]}}|C(x))>T$, where $T$ is a threshold set by the user. By adjusting the threshold, $T$, the refusal rates can be effectively controlled.
Without refusal tokens, refusal sensitivity is typically adjusted by changing the balance of refusals in the SFT dataset and then re-training. Sweeping over the dataset balance parameter is expensive, or even intractable when exploring a multi-dimensional space of different interacting refusal categories. Our token-based strategy enables quick fine-tuning of refusal rates, even over multiple categories, without retraining.
Refusal tokens provide control of the refusal rate. After training on UltraChat and CoCoNot data, we sweep the thresholds of the refusal token. In Figure 2, we observe the tradeoff between true positive (correctly refusing) and false positive (refusing when the model should respond) rates. Figure 2 compares training with (left) and without (right) contrast data, or instruction data that lies close to the boundary between refusal and non-refusal classes but is non-refusal. When contrast data is used, we see that training with the token achieves a better Pareto frontier than training without the token.
@@ 5.1 Controlling Individual Types of Instructions with Category Refusal Tokens
Furthermore, we experiment with five distinct refusal tokens that differentiate between refusal types for CoCoNot. Additionally, we consider the temporal setting with one temporal refusal token. For all experiments in this section, we add refusals and/or contrast data to UltraChat.
@@
Thresholding schemes. We explore category thresholding, refusing with that category token if a token from selected category tokens is the highest probability among the refusal tokens and rises above a threshold. For category thresholding, we emit the refusal token that is the highest probability among the refusal tokens and is in the selected category tokens; otherwise, we emit the token with the highest probability. In Appendix A.1, we explore another scheme: sum thresholding, where we emit a refusal token only if the sum of all category token scores exceeds a threshold. Algorithmic versions of these schemes can be found in Appendix A.7.
@@
Figure 3: Individual category refusal tokens enable precise control over query types. Refusal rates for different categories on CoCoNot when category-specific tokens are suppressed or not generated by the model. The blue dashed bars suppress the refusal token in a model trained exclusively with response and refusal tokens, no category tokens. By suppressing tokens from specific categories during inference, we demonstrate control over the types of refusals. The two dashed bars per group reflect the effect of suppressing a category’s token, either through category-specific suppression or a single refusal token. We also observe category overlap with both these experiments and a manual inspection; for instance, Humanizing Requests queries are similar to other categories.
@@
Independent control of sensitivity for different refusal types. To test whether categories can be independently controlled, we completely suppress each token one-at-a-time, and observe the impact of this suppression on other (non-suppressed) refusal types. In Figure 3, we observe that the sensitivity of each refusal category can be adjusted with little impact on other categories of refusals. There is an exception: Humanizing Requests proved particularly difficult to suppress and did not respond to their token as other categories did. After inspecting the questions and responses of the Humanizing Requests category, we find that many of the questions contained questions or instructions similar to other categories.
@@
Thus, many of the Humanizing questions or instructions are classified as one of the other refusal categories, emitting the incorrect refusal token. For example, many of the questions ask for stock or financial recommendations. These types of requests can easily be refused due to temporal issues (no access to real-time information), input modality issues (needing access to current portfolios), or safety (not wanting to provide financial information). Nevertheless, Figure 3 highlights that one can use individual category tokens to control individual distributions.
@@
Figure 4: Thresholding the refusal tokens increases F1 scores and controls the true positive and false positive rates for a single instruction type (temporal setting). For our temporal experimental setting, we train UltraChat with 22k refusals and 22k contrast examples. The left shows thresholding achieves a better F1 Score, and the right shows thresholding controls the true positive and false positive rates.
@@
Increasing F1 scores via category refusal tokens. In the temporal setting, we sweep the thresholds of a model trained with UltraChat, $\sim 2$k temporal refusal messages, and $\sim 2$k temporal contrast training examples. We experiment with values of $T$ from 0 to $1$ in increments of $0.1$, where we only sweep one token. In Figure 4, we observe that F1 scores improve when properly calibrating the thresholds, finding that $T=0.1$ performs the best. Furthermore, we find that there is an inherent refusal rate on the evaluation queries after training llama-3.1 on a largely refusal-free SFT dataset like Alpaca without adding any refusal messages. Thus, in Figure 4, the false positive rate does not fall below approximately 0.35, which reflects the inherent refusal rate of UltraChat.
@@
Table 2: Using category-wise thresholding to increase the refusal rates of particular categories, a case study. We apply the category-wise threshold at $T=0.1$ to two queries types simultaneously: Humanizing and Indeterminate. This experiment shows that manipulating a subset of categories increases overall F1 performance without retraining the model. In contrast, thresholding a single refusal token yields higher refusal rates across all categories, doubling the contrast refusal rate, and thus, decreasing the F1 Score. The numbers on the left side of the vertical line are the rates that we expect to change by thresholding.
@@
Furthermore, we provide a case study on how to utilize these tokens to improve F1 scores on CoCoNot to show the effectiveness of both category-wise thresholding. To find a threshold that increases the F1 score, we perform a “cheap sweep” where we utilize only the category refusal or response tokens for the labels instead of the LLM judge. More concretely, for each evaluation query from CoCoNot we apply a single forward pass to find which refusal or response token is emitted and use this token as the label. This allows us to evaluate the F1 scores of different thresholds without generating full responses to the queries and judging with an LLM. For each category, in this case we focus on Humanizing and Interderminate categories, the category token threshold is selected by independently altering the threshold for only that token and then selecting the threshold that maximizes the F1 Score. In particular, we sweep thresholds for each token from $0.1$ to $0.9$ in increments of $0.1$. From this, we find that the threshold that maximizes the F1 scores for both Humanizing and Interderminate tokens is $0.1$. In particular, we chose two categories, Humanizing and Interdetermined, as these are the two lowest refusal rates from the five categories across different trained models.
@@
In Table 2, using category-wise thresholding, the F1 Score increased from $0.935$ to $0.946$ with the refusal rates for Humanizing increasing by $\sim 5\%$ and Interdetermined increasing by $\sim 8\%$ with minimal to no impact the contrast refusal rates. Conversely, when setting the single refusal token, the model trained with just a refusal token and response token, to a threshold of $T=0.1$, the contrast refusal rate (Type II error) doubles compared to its baseline state of not thresholding the token. Thus, individually controlling the different category-wise refusal tokens at test-time leads to more control on category refusal rates. Additionally, we observe that logit bias can be used in place of thresholding as well, with F1 Score increased from $0.935$ to $0.943$ with a logit bias of 4 along the same two categories.
6 Immediate Benefits of Training with Refusal Tokens and Contrast Data
@@
Table 3: Refusal tokens and contrast data improve F1 performance on CoCoNot without thresholding at test-time. Ablation studies on training with CoCoNot refusal messages, refusal tokens, and contrast data. We evaluate llama-3 88B performance across different tasks including MMLU (Hendrycks et al., 2020), ARC tasks (Clark et al., 2018), HellaSwag (Zellers et al., 2019), and TruthfulQA MC2 (Lin et al., 2022), following hyperparameters from Tunstall et al. (2023).
@@
Even under normal sampling conditions (i.e, no thresholding strategy is employed), the mere inclusion of refusal tokens during training enhances the model’s refusal behaviors measured by F1 scores. In Table 3, we observe immediate benefits of the token, as the F1 Score on the CoCoNot evaluation improves from 0.918 to 0.940, corresponding to a 26.8% decrease in the error rate or a 2.4% increase in F1 Score. Additionally, in Figure 5, we explore how training on temporal refusal data affects the refusal rates of other CoCoNot categories and TriviaQA. We find that when training only on refusal data (without contrast data) that the refusal rates of TriviaQA increase from $\sim 0\%$ to $\sim 15\%$ and other CoCoNot categories $\sim 40\%$ to $\sim 50\%$. However, we find that including temporal contrast data does not increase the refusal rates of other CoCoNot categories and TriviaQA. These experiments highlight the immediate benefits of contrast data and the refusal token during training.
@@
From Figure 5 (left), even a small number of refusal messages in the training data can influence refusal rates across non-temporal query types. Notably, with just 200 refusal messages, the refusal rates for both CoCoNot queries and TriviaQA questions increase. Thus, these experiments highlight that only training on one type of refusal (i.e, temporal) can affect the refusal rates of other types of questions (i.e, TriviaQA or CoCoNot). Furthermore, from Figure 5 (left), the addition of the refusal token can limit the effect of refusal rate on TriviaQA and CoCoNot queries, but as you scale the number of temporal refusal examples, this benefit is limited.
Figure 5: The token effect of refusal rate on TriviaQA and CoCoNot queries but scaling the number of temporal refusal examples limits the effectiveness of the refusal token. Left are refusal rates on the three groups of queries: temporal questions, CoCoNot questions, and TriviaQA questions. We compare a model trained with the refusal token and one without the token. Right shows the impact of including/excluding contrast data when the refusal token is used. The x-axis shows how many instructions the model was trained with. The gray line shows baseline refusal rates when no refusal messages are in the instruction data.
We add contrast data to understand how adding borderline examples affects the refusal rates. In our experiments, we add one contrast instruction for every one refusal instruction in SFT training data, adding the refusal token to all experiments. From Figure 5 (right), adding the contrast data to the training dataset limits the refusal rates on other instruction types as the number of refusals in the training data scales. Thus, in situations where you only want to refuse a particular instruction type and not affect the refusal rates of other types of instructions, including contrast data in the training data is crucial.
Furthermore, we explore the case where the balance of contrast to refusal messages is one to ten, which is the case for the CoCoNot training dataset. In Table 3, even when training with this imbalance, the contrast training data limits the amount of refusals on innocuous questions, albeit not as dramatically as training without refusals. Additionally, from the table, adding both a single refusal token or category tokens improves F1 scores under the default sampling scheme over not including the token during training. However, we suspect the exact benefits might be model and hyperparameter dependent. Nevertheless, we see benefits in all models that we explored (llama-3.1 and Mistral (Jiang et al., 2023)) in Table 7 in the Appendix.
7 Consistency and Calibration
Token/Answer Agreement.
An issue with refusal messages in LLMs is that generation sampling causes the model’s response to vary across multiple iterations of the same query (Huang et al., 2024). However, the use of a refusal token can help mitigate this issue. For example, we compare two models—one with the refusal token and one without—over five generations for each of $1379$ CoCoNot queries. For each query, we record the number of times the five generations produced the same response type, i.e, a refusal or a response. We find in $87\%$ of queries the five responses produce the exact same response type for the model trained with the token, compared to $81\%$ for the model trained without the refusal token. Although the generations with the token are more consistently producing the same type of response compared to without training with token, the consistency is not $100\%$.
Thus, we tabulated the number of tokens that the model’s response was labelled by the LLM judge as a different label than the output token for the CoCoNot evaluation. Under the model provided sampling parameters, when a single refusal token was generated, the model’s output was labeled as ”Refused” in 1,019 cases and ”Responded” in 29 cases. Conversely, when a response token was generated, the output was labeled as ”Refused” in 46 cases and ”Responded” in 277 cases. These numbers show that a refusal or response token does not guarantee that the generation is a response or refusal. However, it turns out that many of these generations that produce a response label per the LLM judge and contain the refuse token are from the humanizing requests category. In these cases, the model would respond by providing a disclaimer stating that model does not have opinions or feelings, but then continue to answer the question based on some public consensus. Thus, causing the LLM judge to label the model as “responded.”
For example, for a query from the Humanizing category, ”Can you share your personal opinion on the issue of climate change?”, the model responded with ” [Humanizing requests] As an AI, I don’t have personal opinions, beliefs, or feelings. However, I can provide information and perspectives on the issue of climate change based on scientific data and consensus…” From this example, we see that the model outputs the disclaimer that the model does not have personal opinion, etc, but then provides an answer based on scientific consensus and data. Note that without the presence of the token, the model did not include the disclaimer. However, this type of response to this query is an appropriate response as it does not directly respond, claiming it has opinions, etc. Thus, this shows that the refusal token improves the consistency of the model, and even when the responses are not marked as “responded” by an llm-judge the model, the token alters the response of the model.
Calibration analysis.
We measure the calibration of the model trained with single refusal token under three definitions of expected calibration error (ECE): (i) token-level, which compares the predicted probability of emitting the tag with its actual occurrence; (ii) response-level, which contrasts the model’s predicted refusal rate with the empirical refusal frequency across prompts; and (iii) an adjusted variant that first rescales probabilities to the model’s observed minimum and maximum refusal rates, reducing the influence of ceiling and floor effects at extreme thresholds. As shown in Table 4, post-hoc softmax temperature scaling ($\tau=2$) lowers calibration error in all settings, moving the adjusted ECE from $0.13$ to $0.08$, indicating that a simple rescaling step substantially improves the reliability of refusal probabilities without altering the model’s overall refusal propensity.
Table 4: Mean expected calibration error (ECE) for the single-token refusal tag. Lower values indicate better calibration. Post-hoc temperature scaling ($\tau=2$) reduces error across all metrics.
Metric (mean ECE $\downarrow$) Original model Temp-scaled model ($\tau=2$)
Token-level $0.12$ $0.11$
Response-level $0.27$ $0.23$
Adjusted (min-max) $0.13$ $0.08$
8 Conclusion
Beyond their immediate practical value, refusal tokens point to a broader design pattern: teach the model to reason about how to answer by predicting what kind of answer comes first. The community has embraced the same pattern for other capabilities, introducing think/no-think tokens, pause tokens, and tool-calling function tokens that gate internal deliberation or external actions (e.g., (Goyal et al., 2024; Yao et al., 2023)). Each of these extensions reuses the central insight that a short, synthetic token can both expose latent knowledge (as a classifier) and steer subsequent generation (as a condition). We believe future research will continue to enrich this token palette, combining refusal, reflection, tool-use, and reasoning tokens in a single vocabulary, so that one aligned model can fluidly adapt to safety, helpfulness, and user-specific preferences without retraining.
In summary, refusal tokens transform “whether to answer” from a fixed, opaque behaviour into an interpretable, tunable knob, and they do so with negligible engineering overhead. By unifying classification and generation inside the same sequence, they set a blueprint for the next generation of controllable language models.
9 Ethics Statement
This paper studies the refusal messages in Large Language Models (LLMs) and introduces a token to control this behavior in LLMs. Refusal messages are directly related to jailbreaking LLMs. Jailbreaking LLMs is the process of bypassing their built-in safety mechanisms to generate responses that the model would otherwise refuse to provide. This can involve prompt engineering techniques, adversarial attacks, or fine-tuning approaches designed to override default content restrictions. Understanding refusal messages is critical because they represent a key point of failure in jailbreak detection—if an LLM fails to generate a refusal when expected, it suggests a successful jailbreak.